### v2

# Decay and production properties of strange double charm pentaquark

Abstract
In this work we investigate the decay and production properties of the strange double-charm pentaquark $P_{ccs}^{++}$ with strangeness S=1S=-1. Building upon our previous work predicting its $J^{P}=1/2^{-}$ molecular configuration, we employ three-point QCD sum rules to calculate its strong decay widths and estimate its production branching ratios via $\Xi_{bc}^{+}$ baryon decays. The total strong decay width into the $\Xi_{cc}\bar{K}$ and $\Omega_{cc}\pi$ final-state channels is determined as $84.58^{+19.25}_{-18.80}$ MeV. Furthermore, using a rescattering mechanism, we analyze the $\Xi_{bc}^{+}\rightarrow D_{s}^{\ast-}\Xi_{cc}^{++}\rightarrow D^{-}P_{ccs}^{++}$ process and estimate the production branching ratio to be $\mathcal{B}r(\Xi_{bc}^{+}\rightarrow D^{-}P_{ccs}^{++})=(4.32_{-1.47}^{+2.02})\times 10^{-6}$. The relatively narrow width and detectable branching ratio suggest that this pentaquark state could be observed in experiments such as LHCb.
Pentaquark states, exotic states, QCD sum rules
pacs:
12.39.Mk, 12.38.Lg, 14.40.Ev, 14.40.Rt
I Introduction
The study of exotic multiquark states, proposed early in 1964 Gell-Mann:1964ewy ; 1964-Zweig-p- , has become a pivotal frontier in hadronic physics, offering profound insights into the nonperturbative dynamics of quantum chromodynamics (QCD) Nielsen:2009uh ; Chen:2016qju ; Richard:2016eis ; Esposito:2016noz ; Ali:2017jda ; Guo:2017jvc ; Albuquerque:2018jkn ; Liu:2019zoy ; Brambilla:2019esw ; Richard:2019cmi ; Faustov:2021hjs ; Chen:2022asf ; Meng:2022ozq . Since the discovery of the first hidden-charm pentaquarks $P_{c}(4380)$ and $P_{c}(4450)$ by the LHCb Collaboration LHCb:2015yax , significant theoretical and experimental efforts have been devoted to unraveling the nature of these states, which lie beyond the conventional quark model. The recent observation of the double-charm tetraquark $T_{cc}^{+}(3875)$ LHCb:2021vvq ; LHCb:2021auc and the strange-charm tetraquark $T_{c\bar{s}}(2900)$ LHCb:2022lzp further highlights the rich spectrum of exotic hadrons and underscores the potential existence of their double-heavy counterparts, such as double-charm pentaquarks.
Theoretical attempts have been made to study the mass spectrum from both the hadronic molecular picture Yan:2018zdt ; Dong:2021bvy ; Chen:2017vai ; Guo:2017vcf ; Zhu:2019iwm ; Wang:2023eng ; Duan:2024uuf ; Wang:2023aob ; Wang:2023ael ; Sheng:2024hkf and the compact pentaquark picture Chen:2021kad ; Xing:2021yid ; Zhou:2018bkn ; Wang:2018lhz ; Park:2018oib , as well as their electromagnetic properties Ozdem:2022vip ; Ozdem:2024yel ; Zhou:2022gra ; Wang:2023ael ; Sheng:2024hkf ; Zhu:2025abk . In our previous work Yang:2024okq , we systematically investigated the mass spectra of strange double-charm pentaquarks with quark content $ccus\bar{d}$ and strangeness S=1S=-1, employing QCD sum rules for both molecular and compact configurations. Among the predicted configurations, the $J^{P}=1/2^{-}$ molecular pentaquark with $\Xi_{cc}\bar{K}$ structure stands out: its mass (4.20 GeV) lies slightly above the $\Xi_{cc}\bar{K}$ threshold, allowing strong decays only into $\Xi_{cc}\bar{K}$ and $\Omega_{cc}\pi$ channels. This suggests a relatively narrow resonance that could manifest as a discernible peak in experimental invariant mass spectra. The proximity to the threshold suppresses the phase space for strong decays, potentially enhancing its experimental detectability.
Meanwhile, the ongoing experimental quest for doubly heavy $\Xi_{bc}^{+}$ baryons offers a pivotal opportunity to unravel the double-charm pentaquark state. Cabibbo-favored decays of $\Xi_{bc}^{+}$, such as $\Xi_{bc}^{+}\to D_{s}^{*-}\Xi_{cc}^{++}$, could generate the $\Xi_{cc}\bar{K}$ pentaquark through rescattering processes. At the quark level, the diagram for the process $\Xi_{bc}^{+}\rightarrow D^{-}P^{++}_{ccs}$ is shown in the left panel of Fig. 1. The weak decay arises from a Cabibbo-favored weak transition $b\rightarrow c(\bar{c}s)$ along with the creation of a $d\bar{d}$ pair from the strong interaction. This diagram, known as the external W emission diagram, is non-factorizable because the $s\bar{c}$ pair produced in the weak interaction ends up in different final-state hadrons. Thus, long-distance contributions play a significant role in the $\Xi_{bc}^{+}\rightarrow D^{-}P_{ccs}^{++}$ process, where the weakly produced $c\bar{s}$ and uccucc pairs hadronize as $D_{s}^{\ast-}$ and $\Xi_{cc}^{++}$, respectively, followed by a strong rescattering between $D^{-}$ and $P_{ccs}^{++}$ mediated by $\bar{K}^{0}$ meson exchange. The corresponding rescattering process at the hadronic level is shown in the right panel of Fig. 1. The rescattering mechanism for final-state-interaction (FSI) effects has been successfully applied to $D$ meson decays Li:2002pj , $B$ meson decays Cheng:2004ru ; Lu:2005mx , and charm baryon decays Han:2021azw ; Jia:2024pyb . Recently, this mechanism has also been applied to the production of tetraquark states in $B$ meson decays Chen:2020eyu ; Yang:2024coj and of hidden-charm pentaquark states $P_{c}$ in $\Xi_{b}$ decays Hsiao:2024szt . In this work, we apply the rescattering mechanism to the production of the double-charm pentaquark via the $\Xi_{bc}^{+}\rightarrow D^{\ast-}_{s}\Xi_{cc}^{++}\rightarrow D^{-}P^{++}_{ccs}$ process with $\bar{K}^{0}$ exchange.
Figure 1: The production of the strange double-charm pentaquark at the quark level (left) and hadronic level (right).
A precise calculation of the decay width and production branching ratios of the double-charm pentaquark is thus critical to quantify its production rates in such channels and to guide experimental searches at facilities like LHCb and Belle II. In this work, we continue our study by calculating the decay width of the double-charm pentaquark using the QCD sum rule method and estimating its production branching ratios in $\Xi_{bc}$ decays. This paper is organized as follows: In Sec. II, we outline the formalism for computing decay widths within the QCD sum rule approach. Sec. III presents the numerical results for the $J^{P}=1/2^{-}$ pentaquark’s decay properties. Sec. IV discusses its production mechanism via $\Xi_{bc}^{+}$ decays and estimates detectable branching ratios. A brief summary is presented in Sec. V.
II Three Point QCD Sum Rule
Over past several decades, the method of QCD sum rule has been proven to be very powerful to study hadron properties Reinders:1984sr ; Shifman:1978bx ; Colangelo:2000dp ; Narison:2002woh . In this section, we shall study the three-point correlation function of several two-body strong decay process $M\rightarrow X+Y$. For the strong decay process $M\rightarrow X+Y$, the corresponding correlator is written as
$\Pi(p,p^{{}^{\prime}},q)=\int\mathrm{d}^{4}x\mathrm{d}^{4}y\;\mathrm{e}^{\mathrm{i}p^{{}^{\prime}}\cdot x}\mathrm{e}^{\mathrm{i}q\cdot y}\langle 0|T\{J_{X}(x)J_{Y}(y)J^{\dagger}_{M}(0)\}|0\rangle,$ (1)
where $J_{M(X,Y)}$ is the interpolating current for the initial(final) state. In this section, we shall consider the $P_{ccs}\Xi_{cc}\bar{K}$ and $P_{ccs}\Omega_{cc}\pi$ strong decay vertices with $K(\pi)$ off shell. We use the following interpolating currents for $P_{ccs}^{++}$ by considering it as a $\Xi_{cc}\bar{K}$ molecule Yang:2024okq :
$\xi_{1}=[\epsilon_{abc}(c_{a}^{T}\mathcal{C}\gamma_{\mu}c_{b})\gamma_{\mu}\gamma_{5}u_{c}][\bar{d}_{d}\gamma_{5}s_{d}],$ (2)
where $\mathcal{C}$ denote the charge conjugate operator, subscript $a\cdots d$ denote the color index and $u,d,s,c$ denote the up, down, strange, charm quark field, respectively. These current can couple to the $P_{ccs}^{++}$ state with $J^{P}=1/2^{-}$ via
$\langle 0|\xi_{1}|P_{ccs}^{1/2^{-}}\rangle=\lambda^{-}_{P_{ccs}}u(p),$ (3)
in which the value of the coupling constant $\lambda^{-}_{P_{ccs}}$ are determined from the two-point mass sum rules established in Ref. Yang:2024okq :
$\lambda^{-}_{P_{ccs}}=(2.3\pm 0.7)\times 10^{-3}\;\mathrm{GeV}^{6}.$ (4)
The interpolating currents for $\bar{K}$ and $\pi^{+}$ mesons can be constructed as
$J_{\bar{K}}=\mathrm{i}\,\bar{d}_{a}\gamma_{5}s_{a},\quad J_{\pi^{+}}=\mathrm{i}\,\bar{d}_{a}\gamma_{5}u_{a},$ (5)
which can coupling to the meson states via
$\langle 0|J_{\bar{K}}|\bar{K}\rangle=f_{\bar{K}}\frac{m_{K}^{2}}{m_{s}}\equiv\lambda_{K},\quad\langle 0|J_{\pi^{+}}|\pi^{+}\rangle=f_{\pi}\frac{m_{\pi}^{2}}{m_{u}+m_{d}}\equiv\lambda_{\pi}.$ (6)
The interpolating currents for double charm baryons are taken as Zhang:2009iya
$\begin{split}J_{\Xi_{cc}}&=\epsilon_{abc}(c^{T}_{a}C\gamma_{\mu}c_{b})\gamma_{\mu}\gamma_{5}u_{c},\\ J_{\Omega_{cc}}&=\epsilon_{abc}(c^{T}_{a}C\gamma_{\mu}c_{b})\gamma_{\mu}\gamma_{5}s_{c},\end{split}$ (7)
which can coupling to the baryon states via
$\begin{split}\langle 0|J_{\Xi_{cc}}|\Xi_{cc}(p,s)\rangle&=f_{\Xi_{cc}}u(p,s),\\ \langle 0|J_{\Omega_{cc}}|\Omega_{cc}(p,s)\rangle&=f_{\Omega_{cc}}u(p,s).\end{split}$ (8)
The coupling constant $g_{P_{ccs}\Xi_{cc}\bar{K}}$ and $g_{P_{ccs}\Omega_{cc}\pi}$ are defined via the effective Lagrangian Zou:2002yy
$\begin{split}\mathcal{L}_{P_{ccs}\Xi_{cc}\bar{K}}&=g_{P_{ccs}\Xi_{cc}\bar{K}}P_{ccs}\bar{\Xi}_{cc}\bar{K},\\ \mathcal{L}_{P_{ccs}\Omega_{cc}\pi}&=g_{P_{ccs}\Omega_{cc}\pi}P_{ccs}\bar{\Omega}_{cc}\bar{\pi},\end{split}$ (9)
thus the transition matrix element can be obtained as
$\begin{split}\langle\Xi_{cc}(p^{{}^{\prime}},s^{{}^{\prime}})\bar{K}(q)|P_{ccs}(p)\rangle&=g_{P_{ccs}\Xi_{cc}\bar{K}}\bar{u}_{\Xi_{cc}}(p^{{}^{\prime}},s^{{}^{\prime}})u_{P_{ccs}}(p,s),\\ \langle\Omega_{cc}(p^{{}^{\prime}},s^{{}^{\prime}})\pi(q)|P_{ccs}(p)\rangle&=g_{P_{ccs}\Omega_{cc}\pi}\bar{u}_{\Omega_{cc}}(p^{{}^{\prime}},s^{{}^{\prime}})u_{P_{ccs}}(p,s).\end{split}$ (10)
With the above coupling relations and transition matrix element, we can obtain the three-point correlation function Eq. (1) for $P_{ccs}^{++}\rightarrow\Xi_{cc}\bar{K}$ on the phenomenological side
[…]
and for $P_{ccs}^{++}\rightarrow\Omega_{cc}\pi$ process
[…]
On the OPE side, we can evaluate the correlation function with standard QCD sum rule approach. To establish a sum rule for the coupling constant, we will pick out the $1/q^{2}$ terms around the pole $q^{2}\sim 0$ with the structure $\not${p} in the OPE series and then match both sides of the sum rule. To apply sum rules appropriately, we shall calculated at $Q^{2}$ far away from the on-shell mass $m_{K}^{2}$ to ensure the approximation $p^{2}=p^{{}^{\prime}2}=P^{2}$ valid. After performing the Borel transform $P^{2}\rightarrow M_{B}^{2}$ on both phenomenological and OPE sides, we obtain the strong coupling for $P_{ccs}\Xi_{cc}\bar{K}$ vertex
$g_{P_{ccs}\Xi_{cc}\bar{K}}(s_{0},M_{B}^{2})=\frac{1}{\lambda_{P_{ccs}}^{-}\lambda_{\Xi_{cc}}\lambda_{K}(m_{P_{ccs}}+m_{\Xi_{cc}})}\frac{m_{P_{ccs}}^{2}-m_{\Xi_{cc}}^{2}}{\mathrm{e}^{-m_{\Xi_{cc}}^{2}/M_{B}^{2}}-\mathrm{e}^{-m_{P_{ccs}}^{2}/M_{B}^{2}}}\left(\frac{Q^{2}+m_{K}^{2}}{Q^{2}}\right)\left(\int_{s_{<}}^{s_{0}}\mathrm{d}s\;\rho(s)\mathrm{e}^{-s/M_{B}^{2}}+R(M_{B}^{2})\right),$ (13)
and strong coupling for $P_{ccs}\Omega_{cc}\pi$ vertex
$g_{P_{ccs}\Omega_{cc}\pi}(s_{0},M_{B}^{2})=\frac{1}{\lambda_{P_{ccs}}^{-}\lambda_{\Omega_{cc}}\lambda_{\pi}(m_{P_{ccs}}+m_{\Omega_{cc}})}\frac{m_{P_{ccs}}^{2}-m_{\Omega_{cc}}^{2}}{\mathrm{e}^{-m_{\Omega_{cc}}^{2}/M_{B}^{2}}-\mathrm{e}^{-m_{P_{ccs}}^{2}/M_{B}^{2}}}\left(\frac{Q^{2}+m_{\pi}^{2}}{Q^{2}}\right)\left(\int_{s_{<}}^{s_{0}}\mathrm{d}s\;\rho(s)\mathrm{e}^{-s/M_{B}^{2}}+R(M_{B}^{2})\right),$ (14)
where the continuum threshold $s_{0}=22.3\;\mathrm{GeV}^{2}$ is taken from the two-point mass sum rules in Ref. Yang:2024okq .
Using the operator product expansion (OPE) method, the three-point function can also be evaluated at the quark-gluonic level as a function of various QCD parameters. To evaluate the Wilson coefficients, we adopt the quark propagator in momentum space and the propagator
[…]
where $Q$ represents the heavy quark $c$ or $b$, $q$ represents the light quark $u,d,s$, the superscripts $a,b$ denote the color indices. In this work, we will evaluate Wilson coefficients of the correlation function up to dimension nine condensates at the leading order in $\alpha_{s}$. The spectrum function $\rho(s)$ in Eqs. (13), (14) are given in Appendix A. We shall discuss the detail to obtain suitable parameter working regions in QCD sum rule analysis in next section.
III Numerical Analysis
In this section we perform the three-point QCD sum rule analysis for double heavy molecular pentaquark systems using the interpolating currents in Eq. (2). We use the standard values of various QCD condensates as $\langle\bar{q}q\rangle(1\mathrm{GeV})=-(0.24\pm 0.03)^{3}\;\mathrm{GeV}^{3}$, $\langle\bar{q}g_{s}\sigma\cdot Gq\rangle(1\mathrm{GeV})=-M_{0}^{2}\langle\bar{q}q\rangle$, $M_{0}^{2}=(0.8\pm 0.2)\;\mathrm{GeV}^{2}$, $\langle\bar{s}s\rangle/\langle\bar{q}q\rangle=0.8\pm 0.1$, $\langle g_{s}^{2}GG\rangle(1\mathrm{GeV})=(0.48\pm 0.14)\;\mathrm{GeV}^{4}$ at the energy scale $\mu=1$GeV Narison:2002woh ; Jamin:2001zr ; Jamin:1998ra ; Ioffe:1981kw ; Chung:1984gr ; Dosch:1988vv ; Khodjamirian:2011ub ; Francis:2018jyb and $m_{s}(2\;\mathrm{GeV})=95^{+9}_{-3}\;\mathrm{MeV}$, $m_{c}(m_{c})=1.27^{+0.03}_{-0.04}\;\mathrm{GeV}$, $m_{b}(m_{b})=4.18_{-0.03}^{+0.04}\;\mathrm{GeV}$ from the Particle Data GroupParticleDataGroup:2022pth . We also take into account the energy-scale dependence of the above parameters from the renormalization group equation Wang:2025sic
$m_{s}(\mu)=m_{s}(2\mathrm{GeV})\left[\frac{\alpha_{s}(\mu)}{\alpha_{s}(2\mathrm{GeV})}\right]^{\frac{12}{33-2n_{f}}},$
$m_{c}(\mu)=m_{c}(m_{c})\left[\frac{\alpha_{s}(\mu)}{\alpha_{s}(m_{c})}\right]^{\frac{12}{33-2n_{f}}},$
$m_{b}(m_{b})=m_{b}(m_{b})\left[\frac{\alpha_{s}(\mu)}{\alpha_{s}(m_{b})}\right]^{\frac{12}{33-2n_{f}}},$
$\langle\bar{q}q\rangle(\mu)=\langle\bar{q}q\rangle(1\mathrm{GeV})\left[\frac{\alpha_{s}(1\mathrm{GeV})}{\alpha_{s}(\mu)}\right]^{\frac{12}{33-2n_{f}}},$
$\langle\bar{s}s\rangle(\mu)=\langle\bar{s}s\rangle(1\mathrm{GeV})\left[\frac{\alpha_{s}(1\mathrm{GeV})}{\alpha_{s}(\mu)}\right]^{\frac{12}{33-2n_{f}}},$ (17)
$\langle\bar{q}g_{s}\sigma\cdot Gq\rangle(\mu)=\langle\bar{q}g_{s}\sigma\cdot Gq\rangle(1\mathrm{GeV})\left[\frac{\alpha_{s}(1\mathrm{GeV})}{\alpha_{s}(\mu)}\right]^{\frac{2}{33-2n_{f}}},$
$\langle\bar{s}g_{s}\sigma\cdot Gs\rangle(\mu)=\langle\bar{s}g_{s}\sigma\cdot Gs\rangle(1\mathrm{GeV})\left[\frac{\alpha_{s}(1\mathrm{GeV})}{\alpha_{s}(\mu)}\right]^{\frac{2}{33-2n_{f}}},$
$\alpha_{s}(\mu)=\frac{1}{b_{0}t}\left[1-\frac{b_{1}}{b_{0}}\frac{\mathrm{log}t}{t}+\frac{b_{1}^{2}(\mathrm{log}^{2}t-\mathrm{log}t-1)+b_{0}b_{2}}{b_{0}^{4}t^{2}}\right],$
where $t=\mathrm{log}\frac{\mu^{2}}{\Lambda^{2}}$, $b_{0}=\frac{33-2n_{f}}{12\pi}$, $b_{1}=\frac{153-19n_{f}}{24\pi^{2}}$, $b_{2}=\frac{2857-\frac{5033}{9}n_{f}+\frac{325}{27}n_{f}^{2}}{128\pi^{3}}$, $\Lambda=$210 MeV, 292 MeV and 332 MeV for the flavors $n_{f}=$5, 4 and 3, respectively. In this work, we evolve all the input parameters to the energy scale $\mu=m_{c}$ for our sum rule analysis. The parameters for the $K$ and $\pi$ mesons and double charm baryons are adopted in Tab. 1.
Table 1: The values of the hadronic parameters $m_{H}$ and $f_{H}$ in the work taken from Refs. ParticleDataGroup:2022pth ; Brown:2014ena ; Wang:2010hs ; Shi:2019hbf .
Meson(M) Mass $m_{M}$[GeV] Decay constant $f_{M}$[GeV] Baryon(B) Mass $m_{B}$[GeV] Decay constant $f_{B}$[GeV]
$\pi$ 0.140 $0.16\pm 0.04$ $\Xi_{cc}^{++}$ 3.621 0.109
$K$ 0.494 $0.16\pm 0.02$ $\Omega_{cc}^{+}$ 3.738 0.138
III.1 Strong coupling $g_{P_{ccs}\Xi_{cc}\bar{K}}$
Figure 2: The dependence of the strong coupling $g_{P_{ccs}\Xi_{cc}\bar{K}}$ on the Borel mass $M_{B}^{2}$ (left panel) and transfer momentum $Q^{2}$ (right panel). On the left panel, the transfer momentum is set to be $Q^{2}=m_{\Xi_{cc}}^{2}\sim 13.1\;\mathrm{GeV}^{2}$. On the right panel, the red dots denote the value from Eq. (13) with $s_{0}=22.3\;\mathrm{GeV}^{2}$ and $M_{B}^{2}=1.77\;\mathrm{GeV}^{2}$. The blue solid line is the exponential fitting curve. The two dashed blue lines denote the upper and lower boundary of the uncertainty from various condensates, quark masses and hadronic parameters.
In the left panel of Fig. 2, we show the variation of the coupling constant $g_{P_{ccs}\Xi_{cc}\bar{K}}(Q^{2})$ with the Borel mass $M_{B}^{2}$ at $Q^{2}=m_{\Xi_{cc}}^{2}\sim 13.1\;\mathrm{GeV}^{2}$. Such a momentum point is chosen far away from $m_{K}^{2}$ so that it can be safely ignored and the OPE series is valid in this region. We find that the coupling constant $g_{P_{ccs}\Xi_{cc}\bar{K}}(Q^{2})$ has a maximum value at $M_{B}^{2}\sim 1.77\;\mathrm{GeV}^{2}$, around which it has minimal dependence on the non-physical parameter $M_{B}^{2}$. To extrapolate the coupling constant from the valid QCD sum rule working region to the physical pole $Q^{2}=-m_{K}^{2}$, we fit the sum rule result for $s_{0}=22.3\;\mathrm{GeV}^{2}$ and $M_{B}^{2}=1.77\;\mathrm{GeV}^{2}$ with exponential model
$g_{P_{ccs}\Xi_{cc}\bar{K}}(Q^{2})=a\,\mathrm{e}^{-b\,Q^{2}}.$ (18)
The fitting curve is shown in the right panel of Fig. 2, and the result is as follow:
$g_{P_{ccs}\Xi_{cc}\bar{K}}(Q^{2})=-(0.449\pm 0.054\;\mathrm{GeV}^{-3})\mathrm{e}^{-(0.001\pm 0.000)\;\mathrm{GeV}^{-2}\;Q^{2}}.$ (19)
With the on-shell condition $Q^{2}=-m_{K}^{2}$, we can obtain the strong coupling constant $g_{P_{ccs}\Xi_{cc}\bar{K}}(-m_{K}^{2})=-(0.45\pm 0.05)\;\mathrm{GeV}^{2}$. From the matrix element, we can obtain the decay width for $P_{ccs}^{++}\rightarrow\Xi_{cc}^{++}\bar{K}^{0}$ process:
$\Gamma(P_{ccs}^{++}\rightarrow\Xi_{cc}^{++}\bar{K}^{0})=\frac{\sqrt{\lambda(m_{P_{ccs}}^{2},m_{\Xi_{cc}}^{2},m_{K}^{2})}}{8\pi m_{P_{ccs}}^{2}}g_{P_{ccs}\Xi_{cc}\bar{K}}(-m_{K}^{2})\left((m_{P_{ccs}}+m_{\Xi_{cc}})^{2}-m_{K}^{2}\right).$ (20)
Substitute the above on-shell coupling, we can obtain the decay width as
$\Gamma(P_{ccs}^{++}\rightarrow\Xi_{cc}^{++}\bar{K}^{0})=65.02\pm 15.69\,\mathrm{MeV}.$ (21)
III.2 Strong coupling $g_{P_{ccs}\Omega_{cc}\pi}$
Figure 3: The dependence of the strong coupling $g_{P_{ccs}\Omega_{cc}\pi}$ on the Borel mass $M_{B}^{2}$ (left panel) and transfer momentum $Q^{2}$ (right panel ). On the left panel, the transfer momentum is set to be $Q^{2}=m_{\Omega_{cc}}^{2}\sim 13.8\;\mathrm{GeV}^{2}$. On the right panel, the red dots denote the value from Eq. (14) with $s_{0}=22.3\;\mathrm{GeV}^{2}$ and $M_{B}^{2}=1.74\;\mathrm{GeV}^{2}$. The blue solid line is the exponential fitting curve. The two dashed blue lines denote the upper and lower boundary of the uncertainty from various condensates, quark masses and hadronic parameters.
As for $P_{ccs}\Omega_{cc}\pi$ process, in the left panel of Fig. 3, we show the variation of the coupling constant $g_{P_{ccs}\Omega_{cc}\pi}(Q^{2})$ with the Borel mass $M_{B}^{2}$ at $Q^{2}=m_{\Omega_{cc}}^{2}\sim 13.8\;\mathrm{GeV}^{2}$. Such a momentum point is chosen far away from $m_{\pi^{\ast}}^{2}$ so that it can be safely ignored and the OPE series is valid in this region. We find that the coupling constant $g_{P_{ccs}\Omega_{cc}\pi}(Q^{2})$ has a maximum value at $M_{B}^{2}\sim 1.74\;\mathrm{GeV}^{2}$, around which it has minimal dependence on the non-physical parameter $M_{B}^{2}$. We find that the results can be well fitted by the exponential model
$g_{P_{ccs}\Omega_{cc}\pi}(Q^{2})=a\;\mathrm{e}^{-b\;Q^{2}}.$ (22)
The fitting curve is shown in the right panel of Fig. 3, and the result is as follow:
$g_{P_{ccs}\Omega_{cc}\pi}(Q^{2})=-(0.193^{+0.052}_{-0.056}\;\mathrm{GeV}^{-3})\mathrm{e}^{-(0.000\pm 0.00)\;\mathrm{GeV}^{-2}\;Q^{2}}.$ (23)
With the on-shell condition $Q^{2}=-m_{\pi}^{2}$, we can obtain the strong coupling constant $g_{P_{ccs}\Omega_{cc}\pi}(-m_{\pi}^{2})=-0.193$. From the matrix element, we can obtain the decay width for $P_{ccs}^{++}\rightarrow\Xi_{cc}^{++}\bar{K}^{0}$ process:
$\Gamma(P_{ccs}^{++}\rightarrow\Omega_{cc}^{+}\pi^{+})=\frac{\sqrt{\lambda(m_{P_{ccs}}^{2},m_{\Omega_{cc}}^{2},m_{\pi}^{2})}}{8\pi m_{P_{ccs}}^{2}}g_{P_{ccs}\Omega_{cc}\pi}(-m_{\pi}^{2})\left((m_{P_{ccs}}+m_{\Omega_{cc}})^{2}-m_{\pi}^{2}\right).$ (24)
Substitute the above on-shell coupling, we can obtain the decay width as
$\Gamma(P_{ccs}^{++}\rightarrow\Omega_{cc}^{+}\pi^{+})=19.56^{+11.15}_{-10.36}\,\mathrm{MeV}.$ (25)
Thus, we can obtain the total strong decay width as
$\begin{split}\Gamma_{P_{ccs}^{++}}&=\Gamma(P_{ccs}^{++}\rightarrow\Xi_{cc}^{++}\bar{K}^{0})+\Gamma(P_{ccs}^{++}\rightarrow\Omega_{cc}^{+}\pi^{+})\\ &=(84.58^{+19.25}_{-18.80})\,\mathrm{MeV}.\end{split}$ (26)
IV Production via Final-State-Interaction
In the framework of rescattering mechanism, the decay $\Xi_{bc}^{+}\rightarrow P_{ccs}^{++}D^{-}$ can most likely proceed as $\Xi_{bc}^{+}\rightarrow D_{s}^{\ast-}\Xi^{++}_{cc}\rightarrow D^{-}P_{ccs}^{++}$ with $K^{0}$ exchange. Under the factorization approach Wirbel:1985ji ; Bauer:1986bm ; Wang:2017mqp , we can get the decay amplitude of $\Xi_{bc}^{+}\rightarrow D_{s}^{\ast}\Xi_{cc}$:
$\mathcal{A}(\Xi_{bc}^{+}\rightarrow D_{s}^{\ast}\Xi_{cc})=\frac{G_{F}}{\sqrt{2}}V_{cb}V_{cs}a_{1}\epsilon^{\ast\mu}\bar{u}_{\Xi_{cc}}\left(A_{1}\gamma_{\mu}\gamma_{5}+A_{2}\frac{p_{\Xi_{cc},\mu}}{m_{\Xi_{bc}}}\gamma_{5}+B_{1}\gamma_{\mu}+B_{2}\frac{p_{\Xi_{cc},\mu}}{m_{\Xi_{bc}}}\right).$ (27)
The above decay amplitudes in the factorization approach are expressed as
$A_{1}$ $=$ $-\lambda_{D_{s}^{\ast}}\left[g_{1}(m_{D_{s}^{\ast}}^{2})+g_{2}(m_{D_{s}^{\ast}}^{2})\frac{m_{\Xi_{cc}}-m_{\Xi_{bc}}}{m_{\Xi_{bc}}}\right],$ (28)
$A_{2}$ $=$ $-2\lambda_{D_{s}^{\ast}}g_{2}(m_{D_{s}^{\ast}}^{2}),$ (29)
$B_{1}$ $=$ $\lambda_{D_{s}^{\ast}}\left[f_{1}(m_{D_{s}^{\ast}}^{2})-f_{2}(m_{D_{s}^{\ast}}^{2})\frac{m_{\Xi_{cc}}+m_{\Xi_{bc}}}{m_{\Xi_{bc}}}\right],$ (30)
$B_{2}$ $=$ $2\lambda_{D_{s}^{\ast}}f_{2}(m_{D_{s}^{\ast}}^{2}),$ (31)
where $G_{F}$ is the Fermi constant, $V_{ik}$ is the CKM matrix elements, $a_{1}$ is the effective Wilson coefficients obtained by the factorization approach Buchalla:1995vs , and $f_{1,2}$ and $g_{1,2}$ are transition form factors of $\Xi_{bc}^{+}\rightarrow\Xi_{cc}D_{s}^{\ast}$ weak decay process. The above form factors can be parametrized as
$F(Q^{2})=\frac{F(0)}{1-\frac{Q^{2}}{m_{\mathrm{fit}}^{2}}+\delta\left(\frac{Q^{2}}{m_{\mathrm{fit}}^{2}}\right)^{2}},$ (32)
where the parameters $F_{0}$, $m_{\mathrm{fit}}$ and $\delta$ are taken in Ref. Wang:2017mqp and listed in Tab. 2.
Table 2: The values of the parameters $F(0),m_{\mathrm{fit}}$ and $\delta$ for the form factors in Eqs. (28)-(31) for $\Xi_{bc}^{+}\rightarrow D^{\ast}_{s}\Xi_{cc}$ process taken from Ref. Wang:2017mqp .
Form Factor $F(0)$ $m_{\mathrm{fit}}$ $\delta$ Form Factor $F(0)$ $m_{\mathrm{fit}}$ $\delta$
$f_{1}$ 0.550 4.45 0.43 $g_{1}$ 0.530 4.57 0.44
$f_{2}$ -0.230 4.07 0.47 $g_{2}$ -0.043 3.90 0.48
The amplitude for $\Xi_{bc}^{+}\rightarrow D_{s}^{\ast-}\Xi_{cc}^{++}\rightarrow D^{-}P_{ccs}^{++}$ process can be written as:
[…]
where
[…]
The corresponding decay width can be written as
$\Gamma(\Xi_{bc}^{+}\rightarrow P_{ccs}^{++}D^{-})=\frac{\sqrt{\lambda(m_{\Xi_{bc}}^{2},m_{D}^{2},m_{P_{ccs}}^{2})}}{16\pi m_{\Xi_{bc}}^{3}}|\mathcal{A}(\Xi_{bc}^{+}\rightarrow D_{s}^{\ast-}\Xi^{++}_{cc}\rightarrow D^{-}P_{ccs}^{++})|^{2}.$ (35)
It should be noted that in some work of final state interaction formalism Chen:2020eyu ; Hsiao:2024szt , the decay amplitude contains the form factor $F(t,m)=(\Lambda^{2}-m_{K}^{2})/(\Lambda^{2}-t)$ for each strong vertices, which is introduced to compensate the off-shell effect of the exchanged particle at the vertices Gortchakov:1995im . In this work, we can contain the off-shell effect with $Q^{2}$-dependent strong coupling, such as $g_{D_{s}^{\ast}DK}(Q^{2})$ and $g_{P_{ccs}\Xi_{cc}\bar{K}}(Q^{2})$. We take the result of strong coupling $g_{D_{s}^{\ast}DK}(Q^{2})$ with QCD sum rule formalism as follow Yang:2024coj :
$g_{D_{s}^{\ast}DK}(Q^{2})=(2.82^{+1.34}_{-0.82}\,\mathrm{GeV^{-2}})\mathrm{e}^{-(0.22\pm 0.00\,\mathrm{GeV}^{-2})Q^{2}}.$ (36)
With strong coupling Eqs. (19), (36) and Eq. (35), and the mass of $\Xi_{bc}^{+}$ taken from the lattice result Brown:2014ena , the decay width of $\Xi_{bc}^{+}\rightarrow P_{ccs}^{++}D^{-}$ process can be calculated as
$\Gamma(\Xi_{bc}^{+}\rightarrow P_{ccs}^{++}D^{-})=(1.17_{-0.40}^{+0.55})\times 10^{-17}\,\mathrm{GeV}.$ (37)
We take the lattice result of lifetime of $\Xi_{bc}^{+}$ Brown:2014ena , the production branching fraction of $\Xi_{bc}^{+}\rightarrow P_{ccs}^{++}D^{-}$ process can be calculate as
$\mathcal{B}r(\Xi_{bc}^{+}\rightarrow P_{ccs}^{++}D^{-})=(4.32_{-1.47}^{+2.02})\times 10^{-6}.$ (38)
V Summary
Based on our previous calculations of mass spectroscopy Yang:2024okq , we further study the decay and production properties of the exotic strange double charm pentaquark state $P_{ccs}^{++}$ with $J^{P}=1/2^{-}$. We perform the three-point QCD sum rules to calculate the coupling constants of $P_{ccs}^{++}\rightarrow\Xi_{cc}^{++}\bar{K}^{0}$ and $P_{ccs}^{++}\rightarrow\Omega_{cc}^{+}\pi^{+}$ strong decay processes. The partial decay widths of these two process are obtained as
$\begin{split}\Gamma(P_{ccs}^{++}\rightarrow\Xi_{cc}^{++}\bar{K}^{0})&=65.02\pm 15.69\,\mathrm{MeV},\\ \Gamma(P_{ccs}^{++}\rightarrow\Omega_{cc}^{+}\pi^{+})&=19.56^{+11.15}_{-10.36}\,\mathrm{MeV},\end{split}$ (39)
yielding a relative branching ratio $\Gamma(P_{ccs}^{++}\rightarrow\Xi_{cc}^{++}\bar{K}^{0}):\Gamma(P_{ccs}^{++}\rightarrow\Omega_{cc}^{+}\pi^{+})\approx 3.3:1$. The total decay width is predicted as
$\Gamma_{P_{ccs}^{++}}=84.58^{+19.25}_{-18.80}\,\mathrm{MeV}\,.$ (40)
Furthermore, we study the $\Xi_{bc}^{+}\rightarrow D_{s}^{\ast-}\Xi_{cc}^{++}\rightarrow D^{-}P_{ccs}^{++}$ process via the rescattering mechanism to estimate the branching ratio of $\Xi_{bc}^{+}\rightarrow D^{-}P_{ccs}^{++}$ as $(4.32_{-1.47}^{+2.02})\times 10^{-6}$. This value of branching ratio aligns with the productions of hidden-charm pentaquark states in $\Xi_{b}$ decays Wu:2021caw ; Pan:2023hrk ; Wu:2024lud (with a same $b\rightarrow c\bar{c}s$ weak transitions), which typically around $\mathcal{B}r\sim 10^{-6}-10^{-5}$. Combining with the decay branching ratio $\mathcal{B}r(P_{ccs}^{++}\rightarrow\Xi_{cc}^{++}\bar{K}^{0})$, we obtain $\mathcal{B}r(\Xi_{bc}^{+}\rightarrow D^{-}P_{ccs}^{++}\rightarrow D^{-}\bar{K}^{0}\Xi_{cc}^{++})\approx\mathcal{B}r(\Xi_{bc}^{+}\rightarrow D^{-}P_{ccs}^{++})\mathcal{B}r(P_{ccs}^{++}\rightarrow\Xi_{cc}^{++}\bar{K}^{0})=3.3\times 10^{-6}$, providing a benchmark for future experiments.
To date, LHCb has pursued $\Xi_{bc}^{+}$ state in $\Lambda_{c}\pi$, $\Xi_{c}\pi$ and $DKp$ final states without significant signals LHCb:2020iko ; LHCb:2021xba ; LHCb:2022fbu . Nevertheless, high $\Xi_{bc}^{+}$ yields have been expected at future facilities such as MuIC ($\sim 10^{8}$ events/year) Zhao:2025kvq , CEPC/FCC-ee ($\sim 10^{7}$ events/year) Ma:2025ito , LHeC ($\sim 10^{5}$ events/year) Bi:2017nzv , and LHCb Run 3 ($\sim 10^{4}$ events/year) Zhang:2022jst , producing a considerable amount of double charm pentaquark states in the future.
ACKNOWLEDGMENTS
This work is supported by the National Natural Science Foundation of China with Grant Nos. 12175318, the Natural Science Foundation of Guangdong Province of China under Grant No. 2022A1515011922.
Appendix A Spectrum function for three-point correlation of strong vertices
The spectrum function $\rho(s)$ and $R(M_{B}^{2})$ in Eq. (13) is shown as follow
$\rho(s)$ $=$ $\int^{x_{\mathrm{max}}}_{x_{\mathrm{min}}}\mathrm{d}x\int^{y_{\mathrm{max}}}_{y_{\mathrm{min}}}\mathrm{d}y\frac{3}{512\pi^{5}}y\Big(\pi^{2}\langle GG\rangle m_{s}(2\langle\bar{q}q\rangle-\langle\bar{s}s\rangle)((x-2)y+1)-\frac{x}{y-1}\left(\langle GG\rangle+16\pi^{2}m_{s}(\langle\bar{s}s\rangle-2\langle\bar{q}q\rangle)\right)$ (41)
$\Delta(x,y,s)\left(3(x-1)\Delta(x,y,s)+m_{c}^{2}+2s(x-1)(y-1)y\right)\Big),$
$R(M_{B}^{2})$ $=$ $\int^{1}_{0}\mathrm{d}x\int^{1}_{0}\mathrm{d}y\frac{\langle GG\rangle m_{s}(\langle\bar{s}s\rangle-2\langle\bar{q}q\rangle)((x-2)y+1)}{512\pi^{3}(x-1)^{2}(y-1)^{2}}\Big(s_{1}(x-1)^{2}(y-1)^{2}y-m_{c}^{2}(x(y(4((x-1)x+1)y-2x-5)+3)$ (42)
$+y-1)\Big)\mathrm{e}^{-s_{1}/M_{B}^{2}},$
where $x_{\mathrm{min}}=0$, $x_{\mathrm{max}}=\frac{1-2\sqrt{m_{c}^{2}/s}}{(1-\sqrt{m_{c}^{2}/s})^{2}}$, $y_{\mathrm{min}}=\frac{s(1-x)+m_{c}^{2}x-\sqrt{(s(1-x)+m_{c}^{2}x)^{2}-4m_{c}^{2}s(1-x)}}{2s(1-x)}$, $y_{\mathrm{max}}=\frac{s(1-x)+m_{c}^{2}x+\sqrt{(s(1-x)+m_{c}^{2}x)^{2}-4m_{c}^{2}s(1-x)}}{2s(1-x)}$, $\Delta(x,y,s)=-s(1-y)y+\frac{m_{c}^{2}(1-xy)}{1-x}$, $s_{1}=\frac{m_{c}^{2}(1-xy)}{(1-x)(1-y)y}$.
The spectrum function $\rho(s)$ and $R(M_{B}^{2})$ in Eq. (14) is shown as follow
$\rho(s)$ $=$ $\int^{x_{\mathrm{max}}}_{x_{\mathrm{min}}}\mathrm{d}x\int^{y_{\mathrm{max}}}_{y_{\mathrm{min}}}\mathrm{d}y\frac{3}{512\pi^{5}(y-1)}\langle GG\rangle xy\Delta(x,y,s)\left(-3(x-1)\Delta(x,y,s)-m_{c}^{2}-2s(x-1)(y-1)y\right)$ (43)
$+\int^{z_{\mathrm{max}}}_{z_{\mathrm{min}}}\mathrm{d}z\frac{3\langle GG\rangle m_{s}\langle\bar{s}s\rangle(z-1)}{128\pi^{3}}$
$R(M_{B}^{2})$ $=$ $\int^{1}_{0}\mathrm{d}z\frac{\langle GG\rangle m_{s}\langle\bar{s}s\rangle}{256\pi^{3}z}\left(m_{c}^{2}z-s_{2}\left(4z^{2}-6z+2\right)\right)\mathrm{e}^{-s_{2}/M_{B}^{2}}$ (44)
where $z_{\mathrm{min}}=\frac{1}{2}\left(1-\sqrt{1-4m_{c}^{2}/s}\right)$, $z_{\mathrm{max}}=\frac{1}{2}\left(1+\sqrt{1-4m_{c}^{2}/s}\right)$, and $s_{2}=\frac{m_{c}^{2}}{(1-z)z}$.