### v4

# Decay and production properties of strange double charm pentaquark

Abstract
In this work we investigate the decay and production properties of the strange double-charm pentaquark $P_{ccs}^{++}$ with strangeness S=1S=-1. Building upon our previous work predicting its $J^{P}=1/2^{-}$ molecular configuration, we employ three-point QCD sum rules to calculate its strong decay widths and estimate its production branching ratio via $\Xi_{bc}^{+}$ baryon decays. The total strong decay width to the $\Xi_{cc}\bar{K}$ and $\Omega_{cc}\pi$ final-state channels is determined as $85\pm 19$ MeV. Furthermore, using a rescattering mechanism, we analyze the $\Xi_{bc}^{+}\rightarrow D_{s}^{\ast-}\Xi_{cc}^{++}\rightarrow D^{-}P_{ccs}^{++}$ process and estimate the production branching ratio to be $\mathcal{B}r(\Xi_{bc}^{+}\rightarrow D^{-}P_{ccs}^{++})=(4.3_{-1.5}^{+2.0})\times 10^{-6}$. The relatively narrow width and detectable branching ratio suggest the possibility searching for this pentaquark state in the future.
Pentaquark states, exotic states, QCD sum rules
pacs:
12.39.Mk, 12.38.Lg, 14.40.Ev, 14.40.Rt
I Introduction
The study of exotic multiquark states, proposed early in 1964 Gell-Mann:1964ewy ; 1964-Zweig-p- , has become a pivotal frontier in hadronic physics, offering profound insights into the nonperturbative dynamics of quantum chromodynamics (QCD) Nielsen:2009uh ; Chen:2016qju ; Richard:2016eis ; Esposito:2016noz ; Ali:2017jda ; Guo:2017jvc ; Albuquerque:2018jkn ; Liu:2019zoy ; Brambilla:2019esw ; Richard:2019cmi ; Faustov:2021hjs ; Chen:2022asf ; Meng:2022ozq . Since the discovery of the first hidden-charm pentaquarks $P_{c}(4380)$ and $P_{c}(4450)$ by the LHCb Collaboration LHCb:2015yax , significant theoretical and experimental efforts have been devoted to unraveling the nature of these states, which lie beyond the conventional quark model. The recent observation of the double-charm tetraquark $T_{cc}^{+}(3875)$ LHCb:2021vvq ; LHCb:2021auc and the strange-charm tetraquark $T_{c\bar{s}}(2900)$ LHCb:2022lzp further highlights the rich spectrum of exotic hadrons and underscores the potential existence of double-heavy counterparts in the baryon sector, such as double-charm pentaquarks.
Theoretical attempts have been made to study the mass spectrum from both the hadronic molecular picture Yan:2018zdt ; Dong:2021bvy ; Chen:2017vai ; Guo:2017vcf ; Zhu:2019iwm ; Wang:2023eng ; Duan:2024uuf ; Wang:2023aob ; Wang:2023ael ; Sheng:2024hkf and the compact pentaquark picture Chen:2021kad ; Xing:2021yid ; Zhou:2018bkn ; Wang:2018lhz ; Park:2018oib , as well as their electromagnetic properties Ozdem:2022vip ; Ozdem:2024yel ; Zhou:2022gra ; Wang:2023ael ; Sheng:2024hkf ; Zhu:2025abk . In our previous work Yang:2024okq , we systematically investigated the mass spectra of strange double-charm pentaquarks with quark content $ccus\bar{d}$ and strangeness S=1S=-1, employing QCD sum rules for both molecular and compact configurations. Among the predicted configurations, the $J^{P}=1/2^{-}$ molecular pentaquark with $\Xi_{cc}\bar{K}$ structure stands out: its mass (4.20 GeV) lies slightly above the $\Xi_{cc}\bar{K}$ threshold, allowing strong decays only into $\Xi_{cc}\bar{K}$ and $\Omega_{cc}\pi$ channels. This suggests a relatively narrow resonance that could manifest as a discernible peak in experimental invariant mass spectra. The proximity to the threshold suppresses the phase space for strong decays, potentially enhancing its experimental detectability.
Meanwhile, the ongoing experimental quest for doubly heavy $\Xi_{bc}^{+}$ baryons offers a pivotal opportunity to unravel the double-charm pentaquark state. Cabibbo-favored decays of $\Xi_{bc}^{+}$, such as $\Xi_{bc}^{+}\to D_{s}^{*-}\Xi_{cc}^{++}$, could generate the $\Xi_{cc}\bar{K}$ pentaquark through rescattering processes. At the quark level, the diagram for the process $\Xi_{bc}^{+}\rightarrow D^{-}P^{++}_{ccs}$ is shown in the left panel of Fig. 1. The weak decay arises from a Cabibbo-favored weak transition $b\rightarrow c(\bar{c}s)$ along with the creation of a $d\bar{d}$ pair from the strong interaction. This diagram, known as the external W emission diagram, is non-factorizable because the $s\bar{c}$ pair produced in the weak interaction ends up in different final-state hadrons. Thus, long-distance contributions play a significant role in the $\Xi_{bc}^{+}\rightarrow D^{-}P_{ccs}^{++}$ process, where the weakly produced $c\bar{s}$ and uccucc pairs hadronize via $\bar{K}^{0}$ meson exchange interaction, the $D_{s}^{\ast-}$ and $\Xi_{cc}^{++}$ pair converts into a $D^{-}$ and the pentaquark $P_{ccs}^{++}$. The corresponding interaction at the hadronic level is shown in the right panel of Fig. 1. Such a mechanism for final-state-interaction (FSI) effects has been successfully applied to $D$ meson decays Li:2002pj , $B$ meson decays Cheng:2004ru ; Lu:2005mx , and charm baryon decays Han:2021azw ; Jia:2024pyb . Recently, this mechanism has also been applied to the production of tetraquark states in $B$ meson decays Chen:2020eyu ; Yang:2024coj and of hidden-charm pentaquark states $P_{c}$ in $\Xi_{b}$ decays Hsiao:2024szt . In this work, we apply this mechanism to the production of the double-charm pentaquark via the $\Xi_{bc}^{+}\rightarrow D^{\ast-}_{s}\Xi_{cc}^{++}\rightarrow D^{-}P^{++}_{ccs}$ process with $\bar{K}^{0}$ exchange.
Figure 1: The production of the strange double-charm pentaquark at the quark level (left) and hadronic level (right).
A precise calculation of the decay width and production branching ratios of the double-charm pentaquark is thus critical to quantify its production rates in such channels and to guide experimental searches at facilities like LHCb and Belle II. In Ref. Yang:2024okq , we also predicted the masses of possible compact pentaquark states with the same quark content. Notably, two compact configurations were found with masses of $4.25\pm 0.4\;\text{GeV}$, which are close to the mass of the molecular state studied here. It is therefore a natural question how their decay widths and production rates compare with those of the molecular configuration. However, our calculations for these particular compact interpolating currents show that the sum rules analyses for the coupling constants $g_{P_{ccs}\Xi_{cc}\bar{K}}$ and $g_{P_{ccs}\Omega_{cc}\pi}$ did not exhibit a sufficiently stable Borel platform within a reasonable $M_{B}^{2}$ window, leading unreliable predictions of these couplings. Since the hadron decay widths and production rates are highly sensitive to the hadronic couplings, it is difficult to provide quantitative estimations for these compact pentaquark states within the framework of three-point QCD sum rules. In this work, we focus on the well-behaved molecular configuration to study the decay and production properties of strange double charm pentaquark state. We continue our study by calculating the decay width of the double-charm pentaquark using the QCD sum rule method and estimating its production branching ratios in $\Xi_{bc}$ decays. This paper is organized as follows: In Sec. II, we outline the formalism for computing decay widths within the QCD sum rule approach. Sec. III presents the numerical results for the $J^{P}=1/2^{-}$ pentaquark’s decay properties. Sec. IV discusses its production mechanism via $\Xi_{bc}^{+}$ decays and estimates detectable branching ratios. A brief summary is presented in Sec. V.
II Three Point QCD Sum Rule
Over past several decades, the method of QCD sum rule has been proven to be very powerful to study hadron properties Reinders:1984sr ; Shifman:1978bx ; Colangelo:2000dp ; Narison:2002woh . In this section, we shall study the three-point correlation function of several two-body strong decay processes $M\rightarrow X+Y$. For the strong decay process $M\rightarrow X+Y$, the corresponding correlator is written as
$\Pi(p,p^{{}^{\prime}},q)=\int\mathrm{d}^{4}x\mathrm{d}^{4}y\;\mathrm{e}^{\mathrm{i}p^{{}^{\prime}}\cdot x}\mathrm{e}^{\mathrm{i}q\cdot y}\langle 0|T\{J_{X}(x)J_{Y}(y)J^{\dagger}_{M}(0)\}|0\rangle,$ (1)
where $J_{M(X,Y)}$ is the interpolating current for the initial(final) state. In this section, we shall consider the $P_{ccs}\Xi_{cc}\bar{K}$ and $P_{ccs}\Omega_{cc}\pi$ strong decay vertices with $K(\pi)$ off shell. We use the following interpolating current for $P_{ccs}^{++}$ by considering it as a $\Xi_{cc}\bar{K}$ molecule Yang:2024okq :
$\xi_{1}=[\epsilon_{abc}(c_{a}^{T}\mathcal{C}\gamma_{\mu}c_{b})\gamma_{\mu}\gamma_{5}u_{c}][\bar{d}_{d}\gamma_{5}s_{d}],$ (2)
where $\mathcal{C}$ denotes the charge conjugate operator, subscript $a\cdots d$ denotes the color index and $u,d,s,c$ denotes the up, down, strange, charm quark field, respectively. This current can couple to the $P_{ccs}^{++}$ state with $J^{P}=1/2^{-}$ via
$\langle 0|\xi_{1}|P_{ccs}^{1/2^{-}}\rangle=\lambda^{-}_{P_{ccs}}u(p),$ (3)
in which the value of the coupling constant $\lambda^{-}_{P_{ccs}}$ are determined from the two-point mass sum rules established in Ref. Yang:2024okq :
$\lambda^{-}_{P_{ccs}}=(2.3\pm 0.7)\times 10^{-3}\;\mathrm{GeV}^{6}.$ (4)
The interpolating currents for $\bar{K}$ and $\pi^{+}$ mesons can be constructed as
$J_{\bar{K}}=\mathrm{i}\,\bar{d}_{a}\gamma_{5}s_{a},\quad J_{\pi^{+}}=\mathrm{i}\,\bar{d}_{a}\gamma_{5}u_{a},$ (5)
which can couple to the meson states via
$\langle 0|J_{\bar{K}}|\bar{K}\rangle=f_{\bar{K}}\frac{m_{K}^{2}}{m_{s}}\equiv\lambda_{K},\quad\langle 0|J_{\pi^{+}}|\pi^{+}\rangle=f_{\pi}\frac{m_{\pi}^{2}}{m_{u}+m_{d}}\equiv\lambda_{\pi}.$ (6)
The interpolating currents for double charm baryons are taken as Zhang:2009iya
$\begin{split}J_{\Xi_{cc}}&=\epsilon_{abc}(c^{T}_{a}C\gamma_{\mu}c_{b})\gamma_{\mu}\gamma_{5}u_{c},\\ J_{\Omega_{cc}}&=\epsilon_{abc}(c^{T}_{a}C\gamma_{\mu}c_{b})\gamma_{\mu}\gamma_{5}s_{c},\end{split}$ (7)
which can couple to the baryon states via
$\begin{split}\langle 0|J_{\Xi_{cc}}|\Xi_{cc}(p,s)\rangle&=f_{\Xi_{cc}}u(p,s),\\ \langle 0|J_{\Omega_{cc}}|\Omega_{cc}(p,s)\rangle&=f_{\Omega_{cc}}u(p,s).\end{split}$ (8)
The coupling constant $g_{P_{ccs}\Xi_{cc}\bar{K}}$ and $g_{P_{ccs}\Omega_{cc}\pi}$ are defined via the effective Lagrangian Zou:2002yy
$\begin{split}\mathcal{L}_{P_{ccs}\Xi_{cc}\bar{K}}&=g_{P_{ccs}\Xi_{cc}\bar{K}}P_{ccs}\bar{\Xi}_{cc}\bar{K},\\ \mathcal{L}_{P_{ccs}\Omega_{cc}\pi}&=g_{P_{ccs}\Omega_{cc}\pi}P_{ccs}\bar{\Omega}_{cc}\bar{\pi},\end{split}$ (9)
thus the transition matrix element can be obtained as
$\begin{split}\langle\Xi_{cc}(p^{{}^{\prime}},s^{{}^{\prime}})\bar{K}(q)|P_{ccs}(p)\rangle&=g_{P_{ccs}\Xi_{cc}\bar{K}}\bar{u}_{\Xi_{cc}}(p^{{}^{\prime}},s^{{}^{\prime}})u_{P_{ccs}}(p,s),\\ \langle\Omega_{cc}(p^{{}^{\prime}},s^{{}^{\prime}})\pi(q)|P_{ccs}(p)\rangle&=g_{P_{ccs}\Omega_{cc}\pi}\bar{u}_{\Omega_{cc}}(p^{{}^{\prime}},s^{{}^{\prime}})u_{P_{ccs}}(p,s).\end{split}$ (10)
With the above coupling relations and transition matrix element, we can obtain the three-point correlation function Eq. (1) for $P_{ccs}^{++}\rightarrow\Xi_{cc}\bar{K}$ on the phenomenological side
[…]
and for $P_{ccs}^{++}\rightarrow\Omega_{cc}\pi$
[…]
On the OPE side, we can evaluate the correlation function with standard QCD sum rule approach. To establish a sum rule for the coupling constant, we will pick out the $1/Q^{2}$ terms around the pole $Q^{2}\sim 0$ (where $Q^{2}=-q^{2}$) with the structure $\not${p} in the OPE series and then match both sides of the sum rule. The mass of the off-shell meson ($\bar{K}$ or $\pi$) is much smaller than the masses of the baryons ($\Xi_{cc}$, $\Omega_{cc}$, $P_{ccs}$). This justifies the approximation of extracting the $1/q^{2}$ pole term directly in the QCD sum rule analysis, making an exponential extrapolation to the on-shell point unnecessary. After performing the Borel transform $P^{2}\rightarrow M_{B}^{2}$ on both phenomenological and OPE sides, we obtain the strong coupling for the $P_{ccs}\Xi_{cc}\bar{K}$ vertex
$g_{P_{ccs}\Xi_{cc}\bar{K}}(s_{0},M_{B}^{2})=\frac{1}{\lambda_{P_{ccs}}^{-}\lambda_{\Xi_{cc}}\lambda_{K}(m_{P_{ccs}}+m_{\Xi_{cc}})}\frac{m_{P_{ccs}}^{2}-m_{\Xi_{cc}}^{2}}{\mathrm{e}^{-m_{\Xi_{cc}}^{2}/M_{B}^{2}}-\mathrm{e}^{-m_{P_{ccs}}^{2}/M_{B}^{2}}}\left(\int_{s_{<}}^{s_{0}}\mathrm{d}s\;\rho(s)\mathrm{e}^{-s/M_{B}^{2}}+R(M_{B}^{2})\right),$ (13)
and strong coupling for $P_{ccs}\Omega_{cc}\pi$ vertex
$g_{P_{ccs}\Omega_{cc}\pi}(s_{0},M_{B}^{2})=\frac{1}{\lambda_{P_{ccs}}^{-}\lambda_{\Omega_{cc}}\lambda_{\pi}(m_{P_{ccs}}+m_{\Omega_{cc}})}\frac{m_{P_{ccs}}^{2}-m_{\Omega_{cc}}^{2}}{\mathrm{e}^{-m_{\Omega_{cc}}^{2}/M_{B}^{2}}-\mathrm{e}^{-m_{P_{ccs}}^{2}/M_{B}^{2}}}\left(\int_{s_{<}}^{s_{0}}\mathrm{d}s\;\rho(s)\mathrm{e}^{-s/M_{B}^{2}}+R(M_{B}^{2})\right),$ (14)
where the continuum threshold $s_{0}=22.3\;\mathrm{GeV}^{2}$ is taken from the two-point mass sum rules in Ref. Yang:2024okq .
Using the operator product expansion (OPE) method, the three-point function can also be evaluated at the quark-gluonic level as a function of various QCD parameters. To evaluate the Wilson coefficients, we adopt the heavy quark propagator in momentum space and the light quark propagator in coordinate space
[…]
where $Q$ represents the heavy quark $c$ or $b$, $q$ represents the light quark $u,d,s$, the superscripts $a,b$ denote the color indices. In this work, we will evaluate Wilson coefficients of the correlation function up to dimension nine condensates at the leading order in $\alpha_{s}$. The spectral function $\rho(s)$ in Eqs. (13), (14) are given in Appendix A. We shall discuss the detail to obtain suitable parameter working regions in QCD sum rule analysis in next section.
III Numerical Analysis
In this section we perform the three-point QCD sum rule analysis for double heavy molecular pentaquark systems using the interpolating currents in Eq. (2). We use the standard values of various QCD condensates as $\langle\bar{q}q\rangle(1\mathrm{GeV})=-(0.24\pm 0.03)^{3}\;\mathrm{GeV}^{3}$, $\langle\bar{q}g_{s}\sigma\cdot Gq\rangle(1\mathrm{GeV})=-M_{0}^{2}\langle\bar{q}q\rangle$, $M_{0}^{2}=(0.8\pm 0.2)\;\mathrm{GeV}^{2}$, $\langle\bar{s}s\rangle/\langle\bar{q}q\rangle=0.8\pm 0.1$, $\langle g_{s}^{2}GG\rangle(1\mathrm{GeV})=(0.48\pm 0.14)\;\mathrm{GeV}^{4}$ at the energy scale $\mu=1$GeV Narison:2002woh ; Jamin:2001zr ; Jamin:1998ra ; Ioffe:1981kw ; Chung:1984gr ; Dosch:1988vv ; Khodjamirian:2011ub ; Francis:2018jyb and $m_{s}(2\;\mathrm{GeV})=95^{+9}_{-3}\;\mathrm{MeV}$, $m_{c}(m_{c})=1.27^{+0.03}_{-0.04}\;\mathrm{GeV}$, $m_{b}(m_{b})=4.18_{-0.03}^{+0.04}\;\mathrm{GeV}$ from the Particle Data GroupParticleDataGroup:2022pth . We also take into account the energy-scale dependence of the above parameters from the renormalization group equation Wang:2025sic
$m_{s}(\mu)=m_{s}(2\mathrm{GeV})\left[\frac{\alpha_{s}(\mu)}{\alpha_{s}(2\mathrm{GeV})}\right]^{\frac{12}{33-2n_{f}}},$
$m_{c}(\mu)=m_{c}(m_{c})\left[\frac{\alpha_{s}(\mu)}{\alpha_{s}(m_{c})}\right]^{\frac{12}{33-2n_{f}}},$
$m_{b}(m_{b})=m_{b}(m_{b})\left[\frac{\alpha_{s}(\mu)}{\alpha_{s}(m_{b})}\right]^{\frac{12}{33-2n_{f}}},$
$\langle\bar{q}q\rangle(\mu)=\langle\bar{q}q\rangle(1\mathrm{GeV})\left[\frac{\alpha_{s}(1\mathrm{GeV})}{\alpha_{s}(\mu)}\right]^{\frac{12}{33-2n_{f}}},$
$\langle\bar{s}s\rangle(\mu)=\langle\bar{s}s\rangle(1\mathrm{GeV})\left[\frac{\alpha_{s}(1\mathrm{GeV})}{\alpha_{s}(\mu)}\right]^{\frac{12}{33-2n_{f}}},$ (17)
$\langle\bar{q}g_{s}\sigma\cdot Gq\rangle(\mu)=\langle\bar{q}g_{s}\sigma\cdot Gq\rangle(1\mathrm{GeV})\left[\frac{\alpha_{s}(1\mathrm{GeV})}{\alpha_{s}(\mu)}\right]^{\frac{2}{33-2n_{f}}},$
$\langle\bar{s}g_{s}\sigma\cdot Gs\rangle(\mu)=\langle\bar{s}g_{s}\sigma\cdot Gs\rangle(1\mathrm{GeV})\left[\frac{\alpha_{s}(1\mathrm{GeV})}{\alpha_{s}(\mu)}\right]^{\frac{2}{33-2n_{f}}},$
$\alpha_{s}(\mu)=\frac{1}{b_{0}t}\left[1-\frac{b_{1}}{b_{0}}\frac{\mathrm{log}t}{t}+\frac{b_{1}^{2}(\mathrm{log}^{2}t-\mathrm{log}t-1)+b_{0}b_{2}}{b_{0}^{4}t^{2}}\right],$
where $t=\mathrm{log}\frac{\mu^{2}}{\Lambda^{2}}$, $b_{0}=\frac{33-2n_{f}}{12\pi}$, $b_{1}=\frac{153-19n_{f}}{24\pi^{2}}$, $b_{2}=\frac{2857-\frac{5033}{9}n_{f}+\frac{325}{27}n_{f}^{2}}{128\pi^{3}}$, $\Lambda=$210 MeV, 292 MeV and 332 MeV for the flavors $n_{f}=$5, 4 and 3, respectively. In this work, we evolve all the input parameters to the energy scale $\mu=m_{c}$ for our sum rule analysis. The parameters adopted for the $K$ and $\pi$ mesons and double charm baryons are displayed in Tab. 1.
Table 1: The values of the hadronic parameters $m_{H}$ and $f_{H}$ in the work taken from Refs. ParticleDataGroup:2022pth ; Brown:2014ena ; Wang:2010hs ; Shi:2019hbf .
Meson(M) Mass $m_{M}$[GeV] Decay constant $f_{M}$[GeV] Baryon(B) Mass $m_{B}$[GeV] Decay constant $f_{B}$[GeV]
$\pi$ 0.140 $0.16\pm 0.04$ $\Xi_{cc}^{++}$ 3.621 0.109
$K$ 0.494 $0.16\pm 0.02$ $\Omega_{cc}^{+}$ 3.738 0.138
III.1 Strong coupling $g_{P_{ccs}\Xi_{cc}\bar{K}}$
Figure 2: The dependence of the strong coupling $g_{P_{ccs}\Xi_{cc}\bar{K}}$ on the Borel mass $M_{B}^{2}$. The transfer momentum is set to be $Q^{2}=m_{\Xi_{cc}}^{2}\sim 13.1\;\mathrm{GeV}^{2}$.
In Fig. 2, we show the variation of the coupling constant $g_{P_{ccs}\Xi_{cc}\bar{K}}(Q^{2})$ with the Borel mass $M_{B}^{2}$ at $Q^{2}=m_{\Xi_{cc}}^{2}\approx 13.1\;\text{GeV}^{2}$. This momentum point is chosen far away from the kaon pole $m_{K}^{2}$ so that the approximation of extracting the $1/Q^{2}$ term in the QCD sum rule is valid. We find that the coupling constant has a maximum value at $M_{B}^{2}\sim 1.77\;\text{GeV}^{2}$, where it shows minimal dependence on the unphysical Borel parameter. Since the kaon mass is much smaller than the masses of the baryons $\Xi_{cc}$ and $P_{ccs}$, the $Q^{2}$-dependence of the coupling constant is negligible in the range from the sum rule working point $Q^{2}=m_{\Xi_{cc}}^{2}$ to the physical on-shell point $Q^{2}=-m_{K}^{2}$. Therefore, we can directly use the coupling constant evaluated at $Q^{2}=m_{\Xi_{cc}}^{2}$ to compute the decay width without performing an extrapolation. The extracted coupling constant is:
$g_{P_{ccs}\Xi_{cc}\bar{K}}(m_{\Xi_{cc}}^{2})=-(0.45\pm 0.05)\;\text{GeV}^{-3}.$ (18)
This value is then used to compute the decay width. From the matrix element, we can obtain the decay width for $P_{ccs}^{++}\rightarrow\Xi_{cc}^{++}\bar{K}^{0}$ process:
$\Gamma(P_{ccs}^{++}\rightarrow\Xi_{cc}^{++}\bar{K}^{0})=\frac{\sqrt{\lambda(m_{P_{ccs}}^{2},m_{\Xi_{cc}}^{2},m_{K}^{2})}}{8\pi m_{P_{ccs}}^{2}}g_{P_{ccs}\Xi_{cc}\bar{K}}(m_{\Xi_{cc}}^{2})\left((m_{P_{ccs}}+m_{\Xi_{cc}})^{2}-m_{K}^{2}\right).$ (19)
Substituting the above coupling, we can obtain the decay width as
$\Gamma(P_{ccs}^{++}\rightarrow\Xi_{cc}^{++}\bar{K}^{0})=65\pm 16\,\mathrm{MeV}.$ (20)
III.2 Strong coupling $g_{P_{ccs}\Omega_{cc}\pi}$
Figure 3: The dependence of the strong coupling $g_{P_{ccs}\Omega_{cc}\pi}$ on the Borel mass $M_{B}^{2}$. The transfer momentum is set to be $Q^{2}=m_{\Omega_{cc}}^{2}\sim 13.8\;\mathrm{GeV}^{2}$.
For the $P_{ccs}\Omega_{cc}\pi$ vertex, we show the variation of the coupling constant $g_{P_{ccs}\Omega_{cc}\pi}(Q^{2})$ with the Borel mass $M_{B}^{2}$ at $Q^{2}=m_{\Omega_{cc}}^{2}\approx 13.8\;\text{GeV}^{2}$ in Fig. 3. Similar to the discussion above, we extract the $1/Q^{2}$ term in the QCD sum rule since the negligible pion mass. The dependence of the coupling constant on the unphysical Borel parameter achieves minimum around $M_{B}^{2}\sim 1.74\;\text{GeV}^{2}$. Neglecting the small $Q^{2}$-dependence, we compute the coupling constant at $Q^{2}=m_{\Omega_{cc}}^{2}$ as
$g_{P_{ccs}\Omega_{cc}\pi}(m_{\Omega_{cc}}^{2})=-(0.19^{+0.05}_{-0.06})\;\text{GeV}^{-3}.$ (21)
From the matrix element, we obtain the decay width for the $P_{ccs}^{++}\to\Omega_{cc}^{+}\pi^{+}$ process:
$\Gamma(P_{ccs}^{++}\rightarrow\Omega_{cc}^{+}\pi^{+})=\frac{\sqrt{\lambda(m_{P_{ccs}}^{2},m_{\Omega_{cc}}^{2},m_{\pi}^{2})}}{8\pi m_{P_{ccs}}^{2}}g_{P_{ccs}\Omega_{cc}\pi}(m_{\Omega_{cc}}^{2})\left((m_{P_{ccs}}+m_{\Omega_{cc}})^{2}-m_{\pi}^{2}\right).$ (22)
Substituting the above coupling, we can obtain the decay width as
$\Gamma(P_{ccs}^{++}\rightarrow\Omega_{cc}^{+}\pi^{+})=20^{+11}_{-10}\,\mathrm{MeV}.$ (23)
Thus, we can obtain the total strong decay width as
$\begin{split}\Gamma_{P_{ccs}^{++}}&=\Gamma(P_{ccs}^{++}\rightarrow\Xi_{cc}^{++}\bar{K}^{0})+\Gamma(P_{ccs}^{++}\rightarrow\Omega_{cc}^{+}\pi^{+})\\ &=(85\pm 19)\,\mathrm{MeV}.\end{split}$ (24)
IV Production via Final-State-Interaction
In the framework of rescattering mechanism, the decay $\Xi_{bc}^{+}\rightarrow P_{ccs}^{++}D^{-}$ can most likely proceed as $\Xi_{bc}^{+}\rightarrow D_{s}^{\ast-}\Xi^{++}_{cc}\rightarrow D^{-}P_{ccs}^{++}$ with $K^{0}$ exchange. Under the factorization approach Wirbel:1985ji ; Bauer:1986bm ; Wang:2017mqp , we can get the decay amplitude of $\Xi_{bc}^{+}\rightarrow D_{s}^{\ast}\Xi_{cc}$:
$\mathcal{A}(\Xi_{bc}^{+}\rightarrow D_{s}^{\ast}\Xi_{cc})=\frac{G_{F}}{\sqrt{2}}V_{cb}V_{cs}a_{1}\epsilon^{\ast\mu}\bar{u}_{\Xi_{cc}}\left(A_{1}\gamma_{\mu}\gamma_{5}+A_{2}\frac{p_{\Xi_{cc},\mu}}{m_{\Xi_{bc}}}\gamma_{5}+B_{1}\gamma_{\mu}+B_{2}\frac{p_{\Xi_{cc},\mu}}{m_{\Xi_{bc}}}\right).$ (25)
The above decay amplitudes in the factorization approach are expressed as
$A_{1}$ $=$ $-\lambda_{D_{s}^{\ast}}\left[g_{1}(m_{D_{s}^{\ast}}^{2})+g_{2}(m_{D_{s}^{\ast}}^{2})\frac{m_{\Xi_{cc}}-m_{\Xi_{bc}}}{m_{\Xi_{bc}}}\right],$ (26)
$A_{2}$ $=$ $-2\lambda_{D_{s}^{\ast}}g_{2}(m_{D_{s}^{\ast}}^{2}),$ (27)
$B_{1}$ $=$ $\lambda_{D_{s}^{\ast}}\left[f_{1}(m_{D_{s}^{\ast}}^{2})-f_{2}(m_{D_{s}^{\ast}}^{2})\frac{m_{\Xi_{cc}}+m_{\Xi_{bc}}}{m_{\Xi_{bc}}}\right],$ (28)
$B_{2}$ $=$ $2\lambda_{D_{s}^{\ast}}f_{2}(m_{D_{s}^{\ast}}^{2}),$ (29)
where $G_{F}$ is the Fermi constant, $V_{ik}$ is the CKM matrix elements, $a_{1}$ is the effective Wilson coefficients obtained by the factorization approach Buchalla:1995vs , and $f_{1,2}$ and $g_{1,2}$ are transition form factors of $\Xi_{bc}^{+}\rightarrow\Xi_{cc}D_{s}^{\ast}$ weak decay process. The above form factors can be parametrized as
$F(Q^{2})=\frac{F(0)}{1-\frac{Q^{2}}{m_{\mathrm{fit}}^{2}}+\delta\left(\frac{Q^{2}}{m_{\mathrm{fit}}^{2}}\right)^{2}},$ (30)
where the parameters $F_{0}$, $m_{\mathrm{fit}}$ and $\delta$ are taken from Ref. Wang:2017mqp are listed in Tab. 2.
Table 2: The values of the parameters $F(0),m_{\mathrm{fit}}$ and $\delta$ for the form factors in Eqs. (26)-(29) for $\Xi_{bc}^{+}\rightarrow D^{\ast}_{s}\Xi_{cc}$ process taken from Ref. Wang:2017mqp .
Form Factor $F(0)$ $m_{\mathrm{fit}}$ $\delta$ Form Factor $F(0)$ $m_{\mathrm{fit}}$ $\delta$
$f_{1}$ 0.550 4.45 0.43 $g_{1}$ 0.530 4.57 0.44
$f_{2}$ -0.230 4.07 0.47 $g_{2}$ -0.043 3.90 0.48
The amplitude for the $\Xi_{bc}^{+}\rightarrow D_{s}^{\ast-}\Xi_{cc}^{++}\rightarrow D^{-}P_{ccs}^{++}$ process can be written as:
[…]
where
[…]
The corresponding decay width can be written as
$\Gamma(\Xi_{bc}^{+}\rightarrow P_{ccs}^{++}D^{-})=\frac{\sqrt{\lambda(m_{\Xi_{bc}}^{2},m_{D}^{2},m_{P_{ccs}}^{2})}}{16\pi m_{\Xi_{bc}}^{3}}|\mathcal{A}(\Xi_{bc}^{+}\rightarrow D_{s}^{\ast-}\Xi^{++}_{cc}\rightarrow D^{-}P_{ccs}^{++})|^{2}.$ (33)
It should be noted that in some works of the final state interaction formalism Chen:2020eyu ; Hsiao:2024szt , the decay amplitude contains the form factor $F(t,m)=(\Lambda^{2}-m_{K}^{2})/(\Lambda^{2}-t)$ for each strong vertex, which is introduced to compensate the off-shell effect of the exchanged particle at the vertices Gortchakov:1995im . In this work, we can obtain the strong coupling, such $g_{P_{ccs}\Xi_{cc}\bar{K}}$. We take the result of strong coupling $g_{D_{s}^{\ast}DK}(Q^{2})$ with QCD sum rule formalism as follow Yang:2024coj :
$g_{D_{s}^{\ast}DK}(Q^{2})=(2.8^{+1.3}_{-0.8}\,\mathrm{GeV^{-2}})\mathrm{e}^{-(0.2\pm 0.0\,\mathrm{GeV}^{-2})Q^{2}}.$ (34)
With the values of the strong couplings given in Eqs. (18), (34) and the result of Eq. (33), and the mass of $\Xi_{bc}^{+}$ taken from the lattice result Brown:2014ena , the decay width of $\Xi_{bc}^{+}\rightarrow P_{ccs}^{++}D^{-}$ process can be calculated as
$\Gamma(\Xi_{bc}^{+}\rightarrow P_{ccs}^{++}D^{-})=(1.2_{-0.4}^{+0.6})\times 10^{-17}\,\mathrm{GeV}.$ (35)
Taking the lattice result of lifetime of $\Xi_{bc}^{+}$ Brown:2014ena , the production branching fraction of $\Xi_{bc}^{+}\rightarrow P_{ccs}^{++}D^{-}$ process becomes
$\mathcal{B}r(\Xi_{bc}^{+}\rightarrow P_{ccs}^{++}D^{-})=(4.3_{-1.5}^{+2.0})\times 10^{-6}.$ (36)
V Summary
Based on our previous calculations of mass spectroscopy Yang:2024okq , we further study the decay and production properties of the exotic strange double charm pentaquark state $P_{ccs}^{++}$ with $J^{P}=1/2^{-}$. We perform the three-point QCD sum rules to calculate the coupling constants of $P_{ccs}^{++}\rightarrow\Xi_{cc}^{++}\bar{K}^{0}$ and $P_{ccs}^{++}\rightarrow\Omega_{cc}^{+}\pi^{+}$ strong decay processes. The partial decay widths of these two processes are obtained as
$\begin{split}\Gamma(P_{ccs}^{++}\rightarrow\Xi_{cc}^{++}\bar{K}^{0})&=65\pm 16\,\mathrm{MeV},\\ \Gamma(P_{ccs}^{++}\rightarrow\Omega_{cc}^{+}\pi^{+})&=20^{+11}_{-10}\,\mathrm{MeV},\end{split}$ (37)
yielding a relative branching ratio $\Gamma(P_{ccs}^{++}\rightarrow\Xi_{cc}^{++}\bar{K}^{0}):\Gamma(P_{ccs}^{++}\rightarrow\Omega_{cc}^{+}\pi^{+})\approx 3.3:1$. The total decay width is predicted as
$\Gamma_{P_{ccs}^{++}}=85\pm 19\,\mathrm{MeV}\,.$ (38)
Furthermore, we study the $\Xi_{bc}^{+}\rightarrow D_{s}^{\ast-}\Xi_{cc}^{++}\rightarrow D^{-}P_{ccs}^{++}$ process via the rescattering mechanism to estimate the branching ratio of $\Xi_{bc}^{+}\rightarrow D^{-}P_{ccs}^{++}$ as $(4.3_{-1.5}^{+2.0})\times 10^{-6}$. This value of branching ratio aligns with the productions of hidden-charm pentaquark states in $\Xi_{b}$ decays Wu:2021caw ; Pan:2023hrk ; Wu:2024lud (having the same $b\rightarrow c\bar{c}s$ weak transition), which typically lie around $\mathcal{B}r\sim 10^{-6}-10^{-5}$. Combining with the decay branching ratio $\mathcal{B}r(P_{ccs}^{++}\rightarrow\Xi_{cc}^{++}\bar{K}^{0})$, we obtain $\mathcal{B}r(\Xi_{bc}^{+}\rightarrow D^{-}P_{ccs}^{++}\rightarrow D^{-}\bar{K}^{0}\Xi_{cc}^{++})\approx\mathcal{B}r(\Xi_{bc}^{+}\rightarrow D^{-}P_{ccs}^{++})\mathcal{B}r(P_{ccs}^{++}\rightarrow\Xi_{cc}^{++}\bar{K}^{0})=3.3\times 10^{-6}$, providing a benchmark for future experiments.
To date, LHCb has pursued the search of the $\Xi_{bc}^{+}$ state exploring the invariant mass of $\Lambda_{c}\pi$, $\Xi_{c}\pi$ and $DKp$ final states without significant signals LHCb:2020iko ; LHCb:2021xba ; LHCb:2022fbu . Nevertheless, high $\Xi_{bc}^{+}$ yields have been expected at future facilities such as MuIC ($\sim 10^{8}$ events/year) Zhao:2025kvq , CEPC/FCC-ee ($\sim 10^{7}$ events/year) Ma:2025ito , LHeC ($\sim 10^{5}$ events/year) Bi:2017nzv , and LHCb Run 3 ($\sim 10^{4}$ events/year) Zhang:2022jst , which will produce a considerable amount of double charm pentaquark states in the future. Our analysis is specifically applied to the molecular interpretation of the $P_{ccs}^{++}$ state. While compact pentaquark configurations with similar masses are also possible Yang:2024okq , a quantitative analysis of their decay properties within the same QCD sum rule framework proved inconclusive due to the instability of the operator product expansion series for the relevant correlation functions. Future studies with alternative theoretical approaches or improved sum rule constructions may help to clarify the properties of such compact states and their potential distinguishability from molecular candidates in experiments.
ACKNOWLEDGMENTS
This work is supported by the National Natural Science Foundation of China with Grant Nos.  12575153 and 12175318.
Appendix A Spectral function for three-point correlation of strong vertices
The spectral function $\rho(s)$ and $R(M_{B}^{2})$ in Eq. (13) is shown as follow
$\rho(s)$ $=$ $\int^{x_{\mathrm{max}}}_{x_{\mathrm{min}}}\mathrm{d}x\int^{y_{\mathrm{max}}}_{y_{\mathrm{min}}}\mathrm{d}y\frac{3}{512\pi^{5}}y\Big(\pi^{2}\langle GG\rangle m_{s}(2\langle\bar{q}q\rangle-\langle\bar{s}s\rangle)((x-2)y+1)-\frac{x}{y-1}\left(\langle GG\rangle+16\pi^{2}m_{s}(\langle\bar{s}s\rangle-2\langle\bar{q}q\rangle)\right)$ (39)
$\Delta(x,y,s)\left(3(x-1)\Delta(x,y,s)+m_{c}^{2}+2s(x-1)(y-1)y\right)\Big),$
$R(M_{B}^{2})$ $=$ $\int^{1}_{0}\mathrm{d}x\int^{1}_{0}\mathrm{d}y\frac{\langle GG\rangle m_{s}(\langle\bar{s}s\rangle-2\langle\bar{q}q\rangle)((x-2)y+1)}{512\pi^{3}(x-1)^{2}(y-1)^{2}}\Big(s_{1}(x-1)^{2}(y-1)^{2}y-m_{c}^{2}(x(y(4((x-1)x+1)y-2x-5)+3)$ (40)
$+y-1)\Big)\mathrm{e}^{-s_{1}/M_{B}^{2}},$
where $x_{\mathrm{min}}=0$, $x_{\mathrm{max}}=\frac{1-2\sqrt{m_{c}^{2}/s}}{(1-\sqrt{m_{c}^{2}/s})^{2}}$, $y_{\mathrm{min}}=\frac{s(1-x)+m_{c}^{2}x-\sqrt{(s(1-x)+m_{c}^{2}x)^{2}-4m_{c}^{2}s(1-x)}}{2s(1-x)}$, $y_{\mathrm{max}}=\frac{s(1-x)+m_{c}^{2}x+\sqrt{(s(1-x)+m_{c}^{2}x)^{2}-4m_{c}^{2}s(1-x)}}{2s(1-x)}$, $\Delta(x,y,s)=-s(1-y)y+\frac{m_{c}^{2}(1-xy)}{1-x}$, $s_{1}=\frac{m_{c}^{2}(1-xy)}{(1-x)(1-y)y}$.
The spectral function $\rho(s)$ and $R(M_{B}^{2})$ in Eq. (14) is shown as follow
$\rho(s)$ $=$ $\int^{x_{\mathrm{max}}}_{x_{\mathrm{min}}}\mathrm{d}x\int^{y_{\mathrm{max}}}_{y_{\mathrm{min}}}\mathrm{d}y\frac{3}{512\pi^{5}(y-1)}\langle GG\rangle xy\Delta(x,y,s)\left(-3(x-1)\Delta(x,y,s)-m_{c}^{2}-2s(x-1)(y-1)y\right)$ (41)
$+\int^{z_{\mathrm{max}}}_{z_{\mathrm{min}}}\mathrm{d}z\frac{3\langle GG\rangle m_{s}\langle\bar{s}s\rangle(z-1)}{128\pi^{3}}$
$R(M_{B}^{2})$ $=$ $\int^{1}_{0}\mathrm{d}z\frac{\langle GG\rangle m_{s}\langle\bar{s}s\rangle}{256\pi^{3}z}\left(m_{c}^{2}z-s_{2}\left(4z^{2}-6z+2\right)\right)\mathrm{e}^{-s_{2}/M_{B}^{2}}$ (42)
where $z_{\mathrm{min}}=\frac{1}{2}\left(1-\sqrt{1-4m_{c}^{2}/s}\right)$, $z_{\mathrm{max}}=\frac{1}{2}\left(1+\sqrt{1-4m_{c}^{2}/s}\right)$, and $s_{2}=\frac{m_{c}^{2}}{(1-z)z}$.